# The 230 GHz Variability of Numerical Models of Sagittarius A* II. The Physical Origins of the Variability

### v1

@@ Abstract
We explored in Chan et al. (2024) how the ion-electron temperature ratio affects certain numerical models of Sagittarius A* (Sgr A*). Specifically, we studied these effects in magnetic-dominated regions in magnetic-arrested disk (MAD), focusing on the $3$-hour variability at $230$ GHz — $M_{\Delta T}$. In this study, we investigate how variations in electron temperature prescription parameter, $R_{\rm Low}$, influence $M_{\Delta T}$ by analyzing a series of general-relativistic raytracing (GRRT) snapshots. In certain black hole models with a spin $a>0$, we discover that increasing $R_{\rm Low}$ renders the photon ring more optically thick, obscuring the varying accretion flows that contribute to the variability. However, as $R_{\rm Low}$ increases further, MAD flux eruptions become more pronounced, compensating for the decrease in $M_{\Delta T}$. For models with a spin $a<0$, although a higher $R_{\rm Low}$ also increases the optical thickness of the fluid, voids within the optically thick gas fail to cover the entire photon ring. Similarly, flux eruptions are more prominent as $R_{\rm Low}$ increases further, contributing to the observed rise in $M_{\Delta T}$ against $R_{\rm Low}$. For black holes with $a\approx 0$, although the effect of increasing optical depth is still present, their $230$ GHz light curves and hence $M_{\Delta T}$ are insensitive to the changes in $R_{\rm Low}$. Furthermore, we find that the variability of the $230$ GHz light curves at $R_{\rm Low}=1$ correlates with fluctuations in the internal energy of the gas near the black hole, indicating that unusual gas heating may be the source of significant $M_{\Delta T}$ seen in simulations. Our findings highlight potential approaches for refining $M_{\Delta T}$ to better align with observations when modelling Sgr A* or other low-luminosity active galactic nuclei.
@@ 1 Introduction
The galactic-center supermassive black hole Sagittarius A* (Sgr A*) is estimated to accrete matter at a rate of $\sim 10^{-5}$ with respect to its Eddington limit $\dot{M}_{\rm Edd}$ (Yuan et al., 2002). Under such a condition, the accretion disk around Sgr A* can be modeled as an advection-dominated, radiative-inefficient accretion flow (RIAF, Quataert, 2003; Begelman, 2012), where the accreting plasma is optically thin and geometrically thick. While ions are more effectively energized via viscous dissipation (Yuan & Narayan, 2014), electrons are the major channels for radiative cooling (Begelman, 2014). The poor Coulomb coupling between ions and electrons prohibits ions from transferring their energy to electrons and transporting them outwards through radiation. Even if viscous dissipation can heat electrons, they cool much more rapidly than ions (Begelman, 2014). As a result, most of the energy is advected into the black hole. This collisionless nature of the plasma naturally leads to a two-temperature state between the electrons and ions, where $T_{e}$ (electron temperature) $\neq$ $T_{i}$ (ion temperature).
An accurate description of the electron temperature is essential in reproducing the electromagnetic observables of Sgr A*. However, such an accurate treatment requires either two-temperature simulations that separately evolve the electron entropy (Dexter et al., 2020a; Scepi et al., 2022; Grigorian & Dexter, 2024), or first-principles, particle-in-cell plasma simulations in the general-relativistic limit (Ball et al., 2016; Galishnikova et al., 2023), which are both computationally expensive. An ad-hoc approach is to assign $T_{e}$ based on the local $T_{i}$ (Dexter et al., 2010). For instance, Dexter et al. (2009) presented synchrotron emission model of Sgr A* images in which $T_{e}=T_{i}$ everywhere; Mościbrodzka et al. (2009) considered $T_{i}$ to be a fraction $R$ of $T_{e}$ and compared model spectral energy distributions with observations; Mościbrodzka & Falcke (2013) allows $R$ to varies across the domain, so that $R$ in the jet is different from that in the disk; Chan et al. (2015) proposed a prescription function in which $R$ is a step-function of the plasma $\beta$ ($\beta=2P/|B|^{2}$, where $P$ is the gas pressure and $B$ is the magnetic field strength) at some threshold $\beta_{\rm Crit}$ (usually set as $1$); Mościbrodzka et al. (2016) proposed the prescription function
@@
where $R_{\rm High}$ ($R_{\rm Low}$) represent the ion-electron temperature ratios in the weakly (strongly) magnetized regime. This function smooths the sharp transition of the step-function prescription by Chan et al. (2015), and the power-law dependencies on $b$ guarantee that radiations from the strong and weak ion-electron coupling regions are easily distinguishable. Although still ad-hoc, both prescriptions are simple and are physically motivated by the results of particle-in-cell simulations that collisionless plasma preferentially heats the ions for $\beta>1$ (Event Horizon Telescope Collaboration et al., 2019). While this prescription function lacks detailed microphysics of the viscous heating ratio between the ions and the electrons, one can easily tune the free parameters to reproduce a huge amount of spectra and images that can be compared with observations and constrain the parameter spaces.
This simple description was used in Event Horizon Telescope Collaboration et al. (2022) to compare $230$ GHz flux between observations and those generated through general-relativistic magnetohydrodynamics (GRMHD) simulations and general-relativistic raytracing (GRRT) post-processing. None of their models simultaneously pass all of the observational constraints, including the $230$ GHz flux variability, which is defined as the ratio $M_{\Delta T}=\sigma_{\Delta T}/\mu_{\Delta T}$, where $\sigma_{\Delta T}$ is the standard deviation and $\mu_{\Delta T}$ is the mean of the $230$ GHz flux over a time $\Delta T$. For the EHT Collaboration study, $\Delta T=530\,GMc^{-3}=3\,\mathrm{hours}$. The reason why some theoretical models are more variable than observations remains uncertain. In Chan et al. (2024), we try to address this problem (and to our knowledge, this is the first-ever attempt) by varying another free parameter $R_{\rm Low}$ in Equation 1, where it is usually assumed to be $1$. Here, we summarize our findings from Chan et al. (2024)
  1. 1.
@@
However, we have not addressed the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$. In particular
  1. 1.
@@
We aim to examine these questions through an in-depth analysis of the images and simulation snapshots. This paper is structured as follows: We review the procedure of constructing the GRMHD simulation libraries and the GRRT parameter surveys in Section 2, highlighting the parameter spaces we considered in this study. We also describe our process of selecting representative models to address the above questions. In Section 3, we analyze the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$ by examining the images of different $R_{\rm Low}$ and the corresponding GRMHD snapshots. As we will see, the reduction of $M_{\Delta T}$ is an optical depth effect. We discuss the implication of our study in Section 4, and we conclude our studies in Section 5.
@@ 2.1 GRMHD Simulations
We briefly review the procedure of constructing the GRMHD simulation libraries and the parameter surveys for GRRT images of varying $R_{\rm Low}$. The simulation library (named model v3, see, for instance, Table 1 of Event Horizon Telescope Collaboration et al. (2022)) are obtained using open-source code kharma (Prather et al., 2021). kharma solves the ideal GRMHD equations (with $c=1$)
@@
Here, $g_{\mu\nu}$ is the Kerr metric in the funky modified Kerr-Schild coordinates (Wong et al., 2022), $g$ is the determinant of the metric, $\rho$ is the mass density, $\Gamma^{\lambda}_{\nu\kappa}$ is the Christoffel symbol, $B^{i}$ is the magnetic field in the coordinate frame, $u^{\mu}$ is the $4$-velocity, $b^{\nu}$ is the magnetic $4$-vector, and $T^{\mu\nu}$ is the stress-energy tensor (Dhang et al., 2023) defined by
@@
Here $h=1+\gamma/(\gamma-1)P/\rho$ is the specific enthalpy. The $\nabla\cdot\vec{B}=0$ constraint is maintained by the flux-constrained transport scheme (Tóth, 2000). kharma uses the $5$th order Weighted-Essential-Non-Oscillatory scheme (Jiang & Shu, 1996) to reconstruct primitive variables to the cell boundaries and the Lax-Friedrichs solver to construct the Riemann fluxes. kharma evolves the GRMHD equations using the $2$nd order Strong Stability Preserving Runge-Kutta Method (Gottlieb et al., 2011) with a CFL number of $0.7$.
In this work and Chan et al. (2024), we consider only the kharma simulations of the Magnetic Arrested Disc (MAD) state since the MAD models are favoured in Event Horizon Telescope Collaboration et al. (2022). Additionally, recent near-horizon circular polarization measurements and model comparison favour Sgr A* to have dynamically important magnetic fields (Event Horizon Telescope Collaboration et al., 2024). The accreting plasma assumed an ideal gas equation of state with an ion adiabatic index of $4/3$, and the initial conditions being the Fishbone-Moncrief torus (Fishbone & Moncrief, 1976). The torus has an inner radius at $20$ and a pressure maximum at $41$, all in units of $GMc^{-2}$. To initialize the MAD simulation, the initial magnetic field within the torus is set via the vector potential (Wong et al., 2022)
@@
Figure 1: Lowest panel: the $230$ GHz flux light curves for different $R_{\rm Low}$. Here, the black hole spin is $a=+0.94$, $\theta=0.1$ Deg, and $R_{\rm High}=180$. Upper-most panel: Black hole images with $R_{\rm Low}$ correspond to those shown in the lowest panel. In each subplot, we show the $R_{\rm Low}$ and $M_{\Delta T}$ at the upper right corner. The images are taken at the instant marked as squares on top of the light curve. The middle panel is the same as the upper-most panel, but they are taken at the instant marked as triangles. Note that the color scale is normalized to 00 to $1$, and each row of images shares the same color scale. Increasing $R_{\rm Low}$ blocks the photon ring from illuminating, thus reducing the variability. But increasing $R_{\rm Low}$ further would make MAD flux eruptions more visible, hence increasing the variability again.
@@
Figure 2: The time-averaged optical depth of the black hole with $a=+0.94$, $\theta=0.1$ Deg, and $R_{\rm High}=180$ but with varying $R_{\rm Low}$. Here, the optical depth between $0-1$ is in physical units, but those from $1-2$ are mapped via a power-law exponent $\gamma_{\rm map}=\text{log}_{10}(2)/\text{log}_{10}[\rm{MAX}(\tau)]$, so that the maximum $\tau$ across these $4$ figures corresponds to $2$ in the color bar. The photon ring becomes optically thick as $R_{\rm Low}$ increases. The reduced variability is thus an optical depth effect.
@@
Figure 3: The entropy $\mathcal{S}$ (in the log10 scale) contours along the $x-y$ ($x-z$) slice is shown in the upper (lower) panel, demonstrating the small-scale, MAD flux eruptions. This is for the black hole with $a=+0.94$ and is taken at the same instant as the square marker in Figure 1. In the lower panel, we show the $\phi$ coordinate (in Deg) at the lower left- and right-hand corners, respectively. The MAD flux eruption in the image domain has the same morphology as that in the simulation domain, and it exhibits a tube-like structure.
@@
Figure 4: Same as Figure 1, but for the black hole with $a=+0.5$, $\theta=45$ Deg, and $R_{\rm High}=120$. The enhanced variability at the beginning is manifested as a large-scale flux tube generated by MAD flux eruption and is visible as $R_{\rm Low}$ increases.
@@
Figure 5: The magnetic flux tube for the black hole with $a=+0.5$ and $R_{\rm High}=120$ manifested in the $230$ GHz image domain but captured in different viewing angles. The color scale is normalized to $0-1$, and the two figures share the same color scale. We show the $R_{\rm Low}$ and $\theta$ at the upper right corner of each subplot. This is taken at the instant marked as triangles in Figure 7.
@@
Figure 6: The magnetic flux tube shown in Figure 4 and 5 as manifested in the simulation domain. We show the $x-y$ ($x-z$) slice of the density contour (in the log10 scale) with magnetic field lines in the upper (lower) panel. In the lower panel, we show the $\phi$ coordinate (in Deg) at the lower left- and right-hand corners, respectively. This is taken at the same instant as Figure 5.
@@
Figure 8: Same as Figure 7, but with $\theta=90$ Deg, and that we omit the light curves for simplicity. The plasma is still unable to block the photon ring if the black hole is viewed at another angle.
Figure 9: The density (plasma $\beta^{-1}$) contour (both in the log10 scale) along the $x-y$ slice shown in the upper (lower) panel for the black hole with $a=-0.5$. This is to demonstrate the gas voids as manifested in the simulation domain. We overlay velocity streamlines on the density contour. This is taken at the same instant as Figure 8.
@@
Figure 13: Domain-wise modulation index of the gas internal energy $u_{g}$. We show the spin of the black hole at the upper right-hand corner of the subplot. We show in (a) the $x-y$ slice and (b) the $x-z$ slice. The internal energy is most variable close to the event horizon and also along the jet wall, and the variability decreases as $a\rightarrow 0$.

### v2

@@ Abstract
We continue our previous work, Chan et al. (2024), to investigate how variations in the electron temperature prescription parameter, $R_{\rm Low}$, influence the $3$-hour variability at $230$ GHz, $M_{\Delta T}$, in magnetic-arrested disk (MAD) models of Sagittarius A* (Sgr A*), through analyzing a series of general-relativistic magnetohydrodynamics and raytracing simulations. For models with a black hole spin $a>0$, we discovered that increasing $R_{\rm Low}$ renders the photon ring more optically thick, obscuring the varying accretion flows that contribute to the variability. However, as $R_{\rm Low}$ increases further, MAD flux eruptions become more pronounced, compensating for the decrease in $M_{\Delta T}$. For models with a spin $a<0$, although a higher $R_{\rm Low}$ also increases the optical thickness of the fluid, voids within the optically thick gas fail to cover the entire photon ring. Similarly, flux eruptions become more prominent as $R_{\rm Low}$ increases further, contributing to the observed rise in $M_{\Delta T}$ relative to $R_{\rm Low}$. For black holes with a spin $a=0$, although the effect of increasing optical depth is still present, their $230$ GHz light curves, and hence $M_{\Delta T}$, are insensitive to changes in $R_{\rm Low}$. Furthermore, we found that the variability of the $230$ GHz light curves at $R_{\rm Low}=1$ might correlate with fluctuations in the internal energy of the gas near the black hole, and we listed potential causes and solutions to the over-variability problem. Our findings highlight potential approaches for refining $M_{\Delta T}$ to better align with observations when modeling Sgr A*.
@@ 1 Introduction
The galactic-center supermassive black hole Sagittarius A* (Sgr A*) is estimated to accrete matter at a rate of $\sim 10^{-5}$ relative to its Eddington limit, $\dot{M}_{\rm Edd}$ (Yuan et al., 2002). Under such conditions, the accretion disk around Sgr A* can be modeled as an advection-dominated, radiatively inefficient accretion flow (RIAF, Quataert, 2003; Begelman, 2012), where the accreting plasma is optically thin and geometrically thick. While ions are more effectively energized via viscous dissipation (Yuan & Narayan, 2014), electrons are the main channels for radiative cooling (Begelman, 2014). The poor Coulomb coupling between ions and electrons prevents ions from transferring their energy to electrons and transporting it outward through radiation. Even if viscous dissipation can heat electrons, they cool much more rapidly than ions (Begelman, 2014). As a result, most of the energy is advected into the black hole. This collisionless nature of the plasma naturally leads to a two-temperature state between the electrons and ions, where $T_{e}$ (electron temperature) $\neq$ $T_{i}$ (ion temperature).
An accurate description of the electron temperature is essential for reproducing and matching the electromagnetic observables of Sgr A*. However, such an accurate treatment requires either two-temperature simulations that separately evolve the electron entropy (Ressler et al., 2015, 2017; Chael et al., 2018, 2019; Dexter et al., 2020a; Scepi et al., 2022; Grigorian & Dexter, 2024), or first-principles, particle-in-cell plasma simulations in the general-relativistic limit (Ball et al., 2016; Galishnikova et al., 2023), which are both computationally expensive. An ad-hoc approach is to assign $T_{e}$ based on the local $T_{i}$ (Dexter et al., 2010). For instance, Dexter et al. (2009) presented a synchrotron emission model of Sgr A* images in which $T_{e}=T_{i}$ everywhere; Mościbrodzka et al. (2009) considered $T_{i}$ to be a fraction $R$ of $T_{e}$ and compared model spectral energy distributions with observations; Mościbrodzka & Falcke (2013) allows $R$ to vary across the domain, so that $R$ in the jet is different from that in the disk; Chan et al. (2015) proposed a prescription function in which $R$ is a step-function of the plasma $\beta$ ($\beta=2P/b^{2}$, where $P$ is the gas pressure and $b^{2}$ is the square of the magnetic field strength) at some threshold $\beta_{\rm Crit}$ (usually set as $1$); Mościbrodzka et al. (2016) proposed the prescription function:
@@
where $R_{\rm High}$ ($R_{\rm Low}$) represents the ion-electron temperature ratios in the weakly (strongly) magnetized regime. This function smooths the sharp transition of the step-function prescription by Chan et al. (2015), and the power-law dependencies on $b$ guarantee that radiations from the strong and weak ion-electron coupling regions are easily distinguishable. Although still ad-hoc, both prescriptions are simple and are physically motivated by the results of particle-in-cell simulations showing that collisionless plasma preferentially heats the ions for $\beta>1$ (Event Horizon Telescope Collaboration et al., 2019). While this prescription function lacks detailed microphysics of the viscous heating ratio between the ions and the electrons, one can easily tune the free parameters to reproduce a wide range of spectra and images that can be compared with observations and constrain the parameter spaces. Note that the electron temperature is given as:
$T_{e}=\frac{2m_{p}\epsilon}{3k(2+R)}$ (2)
where $\epsilon$ is the specific internal energy of the bulk flow, and we assume that the plasma is an electron-proton plasma such that their internal energies add up to the bulk internal energy. Note that we also assume the electron (proton) is relativistic (non-relativistic) with an adiabatic index of $4/3$ ($5/3$).
This simple description was used in Event Horizon Telescope Collaboration et al. (2022) to compare the $230$ GHz flux between observations and those generated through general-relativistic magnetohydrodynamics (GRMHD) simulations and general-relativistic ray-tracing (GRRT) post-processing. None of their models simultaneously pass all of the observational constraints, including the $230$ GHz flux variability, which is defined as the ratio $M_{\Delta T}=\sigma_{\Delta T}/\mu_{\Delta T}$, where $\sigma_{\Delta T}$ is the standard deviation and $\mu_{\Delta T}$ is the mean of the $230$ GHz flux over a time $\Delta T$. For the EHT Collaboration study, $\Delta T=530\,GMc^{-3}=3\,\mathrm{hours}$. The reason why some theoretical models are more variable than observations remains uncertain. In Chan et al. (2024), we try to address this problem by varying another free parameter $R_{\rm Low}$ in Equation 1, where it is usually assumed to be $1$. Here, we summarize our findings from Chan et al. (2024):
@@
and we refer readers to Figure 1 for a recap on the parameter dependence of $M_{\Delta T}$ against $R_{\rm Low}$. We also remind readers here that the photon ring of a black hole image consists of an infinite series of sub-rings that characterize the number of orbits that a photon completes around the black hole before reaching the observer. Although models with $R_{\rm Low}\neq 1$ have been briefly explored in the contexts of M87 and Sgr A* (Event Horizon Telescope Collaboration et al., 2021, 2022, 2023, 2025), a systematic parameter search and analysis on a wide range of $R_{\rm Low}$ has only been conducted in our previous work Chan et al. (2024), focusing on Sgr A*. However, we have not addressed the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$. In particular:
@@
and we aim to examine these questions through an in-depth analysis of the images and simulation snapshots. This paper is structured as follows: We review the procedure of constructing the GRMHD simulation libraries and the GRRT parameter surveys in Section 2, highlighting the parameter spaces we considered in this study. We also describe our process of selecting representative models to address the above questions. In Section 3, we analyze the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$ by examining the images of different $R_{\rm Low}$ and the corresponding GRMHD snapshots. As we will see, the reduction of $M_{\Delta T}$ is an optical depth effect. We discuss the implications of our study in Section 4, and we conclude our studies in Section 5.
@@ 2.1 GRMHD Simulations
We briefly review the procedure of constructing the GRMHD simulation libraries and the parameter surveys for GRRT images of varying $R_{\rm Low}$. The simulation library (named model v3, see, for instance, Table 1 of Event Horizon Telescope Collaboration et al. (2022)) is obtained using the open-source code kharma (Prather et al., 2021). kharma solves the ideal GRMHD equations (with $c=1$):
@@
and here, $g_{\mu\nu}$ is the Kerr metric in the funky modified Kerr-Schild coordinates (Wong et al., 2022), $g$ is the determinant of the metric, $\rho$ is the mass density, $\Gamma^{\lambda}_{\nu\kappa}$ is the Christoffel symbol, $B^{i}$ is the magnetic field in the coordinate frame, $u^{\mu}$ is the $4$-velocity, $b^{\nu}$ is the magnetic $4$-vector, and $T^{\mu\nu}$ is the stress-energy tensor (Dhang et al., 2023) defined by:
@@
so that $h=1+\gamma/(\gamma-1)P/\rho$ is the specific enthalpy. The $\nabla\cdot\vec{B}=0$ constraint is maintained by the flux-constrained transport scheme (Tóth, 2000). kharma uses the 5th order Weighted-Essential-Non-Oscillatory scheme (Jiang & Shu, 1996) to reconstruct primitive variables to the cell boundaries and the Lax-Friedrichs solver to construct the Riemann fluxes. kharma evolves the GRMHD equations using the 2nd order Strong Stability Preserving Runge-Kutta Method (Gottlieb et al., 2001) with a Courant–Friedrichs–Lewy number of $0.7$.
In this work and Chan et al. (2024), we consider only the kharma simulations of the Magnetic Arrested Disc (MAD) state since the MAD models are favored in Event Horizon Telescope Collaboration et al. (2022). Additionally, recent near-horizon circular polarization measurements and model comparisons favor Sgr A* to have dynamically important magnetic fields (Event Horizon Telescope Collaboration et al., 2024). The accreting plasma assumed an ideal gas equation of state with an ion adiabatic index of $4/3$, and the initial conditions were based on the Fishbone-Moncrief torus (Fishbone & Moncrief, 1976). The torus has an inner radius at $20$ and a pressure maximum at $41$, all in units of $GMc^{-2}$. To initialize the MAD simulation, the initial magnetic field within the torus is set via the vector potential (Wong et al., 2022):
@@
Figure 1: Recapping the parameter dependence of $M_{\Delta T}$ against $R_{\rm Low}$ from Chan et al. (2024). We label the model $\theta$ and $a$ in the upper right-hand corner of each subplot.
@@
Figure 2: Lowest panel: the $230$ GHz flux light curves for different $R_{\rm Low}$. Here, the black hole spin is $a=+0.94$, $\theta=0.1$ Deg, and $R_{\rm High}=180$. Upper-most panel: Black hole images with $R_{\rm Low}$ correspond to those shown in the lowest panel. In each subplot, we show the $R_{\rm Low}$ and $M_{\Delta T}$ in the upper right corner. The images are taken at the instants marked as squares on top of the light curve. The middle panel is the same as the upper-most panel, but the images are taken at the instants marked as triangles. Images are shown in units of brightness temperature $T_{b}$ (in $10^{9}$ K), and all images share the same color scale. Increasing $R_{\rm Low}$ blocks the photon ring from illuminating, thus reducing the variability. However, increasing $R_{\rm Low}$ further makes MAD flux eruptions more visible, hence increasing the variability again.
@@
Figure 3: The time-averaged optical depth of the black hole with $a=+0.94$, $\theta=0.1$ Deg, and $R_{\rm High}=180$, but with varying $R_{\rm Low}$. Here, the optical depth between $0-1$ is in physical units, but those from $1-2$ are mapped via a power-law exponent $\gamma_{\rm map}=\text{log}_{10}(2)/\text{log}_{10}[\rm{MAX}(\tau)]$, so that the maximum $\tau$ across these four figures corresponds to $2$ in the color bar. The photon ring becomes optically thick as $R_{\rm Low}$ increases. The reduced variability is thus an optical depth effect.
@@
Figure 4: The entropy $\mathcal{S}$ (in the log10 scale) contours along the $x-y$ ($x-z$) slice are shown in the upper (lower) panel, demonstrating the small-scale, MAD flux eruptions. This is for the black hole with $a=+0.94$ and is taken at the same instant as the triangular marker in Figure 2. In the lower panel, we show the $\phi$ coordinate (in Deg) at the upper left- and right-hand corners, respectively. The MAD flux eruption in the image domain has the same morphology as that in the simulation domain, and it exhibits a tube-like structure.
@@
Figure 5: Same as Figure 2, but for the black hole with $a=+0.5$, $\theta=45$ Deg, and $R_{\rm High}=120$. The enhanced variability at the beginning of this time chunk is manifested as a large-scale flux tube generated by the MAD flux eruption and becomes more visible as $R_{\rm Low}$ increases.
@@
Figure 6: The magnetic flux tube for the black hole with $a=+0.5$ and $R_{\rm High}=120$ manifested in the $230$ GHz image domain but captured from different viewing angles. Both images are shown in units of brightness temperature $T_{b}$ (in $10^{9}$ K). We show the $R_{\rm Low}$ and $\theta$ in the upper right corner of each subplot. This is taken at the instant marked as triangles in Figure 8.
@@
Figure 7: The magnetic flux tube shown in Figures 5 and 6 as manifested in the simulation domain. We show the $x-y$ ($x-z$) slice of the density contour (in the log10 scale) with magnetic field lines in the upper (lower) panel. In the lower panel, we show the $\phi$ coordinate (in Deg) at the lower left- and right-hand corners, respectively. This is taken at the same instant as Figure 6.
@@
Figure 9: Same as Figure 8, but with $\theta=90$ Deg, and we omit the light curves for simplicity. The plasma is still unable to block the photon ring if the black hole is viewed from another angle.
Figure 10: Same as Figure 7, but for the black hole with $a=-0.5$. This is to demonstrate the gas voids and flux eruptions as manifested in the simulation domain. We omitted vector streamlines for simplicity. This is taken at the same instant as Figure 9.
@@
Figure 14: Transformed light curves for both $R_{\rm Low}=1,10$ are shown for models with (a) $a=+0.94$ and (b) $a=+0.5$. The solid line represents the sample mean, while the shaded region indicates the sample range. Increasing $R_{\rm Low}$ leads to an overall ‘flattening’ of the transformed light curves, indicating a reduced variability.
Figure 15: Computational domain-wise modulation index of the gas internal energy ($M_{3}(u_{g})$), binned emission ($M_{3}(I_{\gamma})$), and the time-averaged binned emission ($\langle I_{\gamma}\rangle$, in the log10 scale). Here, $\langle I_{\gamma}\rangle$ is in arbitrary unit. We show models of $a=+0.94$ (a) and $a=+0.5$ (b). The origin of emission is predominantly along the jet funnel wall and is also variable at the same position. The gas internal energy is also most variable along the jet funnel wall.